The depth of the fluid is 10 m. What is the specific gravity if the pressure of the fluid at 10 m is 100 kPa?

1
Expert's answer
2022-02-09T13:40:30-0500

Answer

Pressure at h=10m is P=100kPa

So "P=dgh\\"

Density of fluid

"d=\frac{P}{gh}\\=\frac{10000}{10\times9.81}\\=101.94kg/m^3"

So specific gravity

"S=\frac{fluid \space density}{water \space density}\\=\frac{101.94}{997}\\=0.102"